How much low fat milk (1% fat) and how much whole milk (4% fat) should be mixed to make 5 gallons of reduced fat milk (2% fat)? Please solve with the systems of linear equation.
Work done so far:
low fat milk= x
whole milk= y
reduced fat= t
0.01x+0.04y= 5x+0.02t

Answer by Alan3354(69443)

How much low fat milk (1% fat) and how much whole milk (4% fat) should be mixed to make 5 gallons of reduced fat milk (2% fat)? Please solve with the systems of linear equation.
Work done so far:
low fat milk= x
whole milk= y
reduced fat= t
0.01x+0.04y= 5x+0.02t
----------------
Close
0.01x + 0.04y = 0.02*5 (total fat)
x + y = 5 (total milk)